OpenStudy (anonymous):

Integral challenge\[\int_{0}^{\infty} \sin (x^2) \ \text{d}x = ?\]

OpenStudy (rational):

.

OpenStudy (usukidoll):

u substitution?

OpenStudy (usukidoll):

like let u = x^2 du = 2dx \[\frac{du}{2} = dx\] and then -cos x because the derivative of cos x is -sin x and if we have that extra negative we can multiply . A negative x negative is positive -(-sinx) = sinx

OpenStudy (usukidoll):

sorry I'm typing this while dealing with a stomachache

OpenStudy (usukidoll):

I just typo'ed dropped the x :(

OpenStudy (usukidoll):

u = x^2 du = 2x dx \[\frac{du}{2x} = dx\]

OpenStudy (usukidoll):

so to gain sin(u) back I need -cos(u) because the derivative of cos(u) is -sin(u) and with that extra negative in place... a negative x a negative is a positive so -(-sin(u)) ->sin (u) \[\frac{-1}{2x}\cos(x^2)+C\]

OpenStudy (usukidoll):

FORGET IT! It looks innocent, yet when I try to do it, it's all screwed up.

OpenStudy (usukidoll):

nugh! there has got to be a way to get rid of that second part!!!

OpenStudy (usukidoll):

first part is fine... second is leave cos(x^2) deal with that fraction which is quotient rule

OpenStudy (michele_laino):

hint: we can write this: \[\large \int_0^{ + \infty } {\sin \left( {{x^2}} \right)} \;dx = \frac{1}{2}\int_{ - \infty }^{ + \infty } {\sin } \left( {{x^2}} \right)\;dx\]

OpenStudy (michele_laino):

now I change variable: \[\large {x^2} = z\] where z is the new integration variable

OpenStudy (anonymous):

There is no anti-derivative for \(\sin x^2\) in terms of elementary functions. But there is a way to solve this integral with elementary functions. Also a little bit of Laplace transforms needed here and first thing to do is changing it to a double integral.

OpenStudy (michele_laino):

\[\large \int_0^{ + \infty } {\sin \left( {{x^2}} \right)} \;dx = \int_0^{ + \infty } {\sin \left( z \right)} \;\frac{{dz}}{{2\sqrt z }}\]

OpenStudy (anonymous):

In case you don't wanna use complex analysis. Michele is on the right track till now.

OpenStudy (michele_laino):

the subsequent function: \[\large \frac{{\sin \left( z \right)}}{{2\sqrt z }}\] is holomorphic in all z-plane except z=0

OpenStudy (anonymous):

you wanna use complex analysis?

OpenStudy (michele_laino):

may I apply complex analysis?

OpenStudy (anonymous):

oh yeah, no problem

OpenStudy (michele_laino):

ok! Then the line of integration is: